OpenStudy (anonymous):

Pleeeeaase Pleeeeaaase help me?? Algebra 2... I've missed a few lessons and am completely lost :"(

jimthompson5910 (jim_thompson5910):

What's your question

OpenStudy (anonymous):

Show all work in dividing \[\frac{ 4t^2-1 }{ 15t+10 }\div \frac{ 4t^2-4t+1 }{ 3t^2+5t+2 }\]

OpenStudy (anonymous):

use complete sentences to explain your process

jimthompson5910 (jim_thompson5910):

were you able to get anywhere?

OpenStudy (anonymous):

@jim_thompson5910 I am completely confuzzled :(

OpenStudy (anonymous):

@Luis_Rivera I am desperate :(

jimthompson5910 (jim_thompson5910):

the first step is to flip the second fraction and multiply so you go from \[\large \frac{ 4t^2-1 }{ 15t+10 }\div \frac{ 4t^2-4t+1 }{ 3t^2+5t+2 }\] to \[\large \frac{ 4t^2-1 }{ 15t+10 }\times \frac{ 3t^2+5t+2 }{ 4t^2-4t+1 }\]

jimthompson5910 (jim_thompson5910):

Now combine the numerators and denominators to get \[\large \frac{ (4t^2-1)(3t^2+5t+2) }{ (15t+10)(4t^2-4t+1) }\] From here, you need to factor each expression as much as possible. Then cancel out the common terms and simplify.

jimthompson5910 (jim_thompson5910):

so what do you get when you factor 4t^2 - 1

jimthompson5910 (jim_thompson5910):

any ideas P.nut1996?

OpenStudy (anonymous):

I don't remember how to factor:(

jimthompson5910 (jim_thompson5910):

use the difference of squares rule

jimthompson5910 (jim_thompson5910):

a^2 - b^2 = (a+b)(a-b)

jimthompson5910 (jim_thompson5910):

2t^2 - 1 2t^2 - 1^2 so how would you use the rule above to factor?

jimthompson5910 (jim_thompson5910):

if you're not sure, then a = 2t and b = 1 now just plug them into a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

Thank you so much for the help!! <3

jimthompson5910 (jim_thompson5910):

that's not the full/final answer of course since you still need to factor and reduce

jimthompson5910 (jim_thompson5910):

but you're welcome